A Complete Bibliography of ACM Transactions on Recommender Systems

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

20 October 2023 Version 1.02

Title word cross-reference

n [7].

ACM [1]. Aggregator [8]. Algorithm [6]. armed [2]. Auditing [6].

Bandits [2]. based [8]. Bottleneck [5]. Bubbles [6].

Challenges [3]. Cold [2]. Cold-start [2]. Collaborative [9]. Completion [15]. Consistency [15].

Data [14]. Debiased [5]. Decision [4]. Decision-Making [4]. Directions [3].

Editorial [1]. Estimating [7]. Evaluating [7]. Experience [2]. Explicitly [8].

Fairness [15]. Features [8]. Filter [6]. Filtering [9]. First [2]. Framework [9].

GCN [8]. Graph [3]. Grocery [10]. Guarantees [15]. Guide [2].

Inaugural [1]. Information [5]. Intent [13]. Intent-Satisfaction [13]. Interaction [14]. Issue [1]. Item [10].

Learning [11, 5].

Making [4]. Methods [3]. Misinformation [6]. Modeling [13]. Multi [2, 12]. Multi-armed [2].

Multi-organization [12]. Music [13].

Networks [3]. Neural [3]. Next [10].

organization [12].

Period [10]. Personalization [14]. Pessimistic [4]. Popularity [8]. Predictions [7]. Problem [2]. Provable [15]. Purchase [10].

Rating [7]. Recommendation [11, 8, 10, 5, 6, 12]. Recommendations [7, 2]. Recommender [1, 7, 3, 4, 15]. Reducing [14]. Reinforcement [11]. Representation [5]. Reproducibility [11]. Reverse [10].

Satisfaction [13]. Self [9]. Self-supervised [9]. SelfCF [9]. Shopping [10]. Simple [9]. start [2]. Streaming [13]. Study [11]. supervised [9]. Survey [3]. Systematic [11]. Systems [11, 7, 3, 4, 15, 1].

Targeted [12]. Temporal [8]. Tensor [15]. Top [7]. Top- [7]. Training [12]. Transactions [1].

Uncertainty [7]. User [14, 2].

via [5]. Video [13].

Weighted [8]. Who [10]. Will [10].

YouTube [6].

References

Chen:2023:ATR

[1] Li Chen and Dietmar Jannach. ACM Transactions on Recommender Systems: Inaugural issue editorial. ACM Transactions on Recommender Systems (TORS), 1(1):1:1–1:??, March 2023. ISSN 2770-

6699. URL https://dl.acm.org/doi/
10.1145/3569454.

Silva:2023:UCS

[2] Nicollas Silva, Thiago Silva, Heitor Werneck, Leonardo Rocha, and Adriano Pereira. User cold-start problem in multiarmed bandits: When the first recommendations guide the user’s experience. ACM Transactions on Recommender Systems (TORS), 1(1):2:1–2:??, March 2023. ISSN 2770-6699. URL https://dl.acm.

org/doi/10.1145/3554819.

Gao:2023:SGN

[3] Chen Gao, Yu Zheng, Nian Li, Yinfeng Li, Yingrong Qin, Jinghua Piao, Yuhan Quan, Jianxin Chang, Depeng Jin, Xiangnan He, and Yong Li. A survey of graph neural networks for recommender systems: Challenges, methods, and directions. ACM Transactions on Recommender Systems (TORS), 1(1):3:1–3:??, March 2023. ISSN 2770-

6699. URL https://dl.acm.org/doi/
10.1145/3568022.

Jeunen:2023:PDM

[4] Olivier Jeunen and Bart Goethals. Pessimistic decision-making for recommender systems. ACM Transactions on Recommender Systems (TORS), 1 (1):4:1–4:??, March 2023. ISSN 2770-

6699. URL https://dl.acm.org/doi/
10.1145/3568029.

Liu:2023:DRL

[5] Dugang Liu, Pengxiang Cheng, Hong Zhu, Zhenhua Dong, Xiuqiang He, Weike Pan, and Zhong Ming. Debiased representation learning in recommendation via information bottleneck. ACM Transactions on Recommender Systems (TORS), 1(1):5:1–5:??, March 2023. ISSN 2770-

6699. URL https://dl.acm.org/doi/
10.1145/3568030.

Srba:2023:AYR

[6] Ivan Srba, Robert Moro, Matus Tomlein, Branislav Pecher, Jakub Simko, Elena Stefancova, Michal Kompan, Andrea Hrckova, Juraj Podrouzek, Adrian Gavornik, and Maria Bielikova. Auditing YouTube’s recommendation algorithm for misinformation filter bubbles. ACM Transactions on Recommender Systems (TORS), 1(1):6:1–6:??, March 2023. ISSN 2770-6699. URL https://dl.acm.

org/doi/10.1145/3568392.

Coscrato:2023:EEU

[7] Victor Coscrato and Derek Bridge. Estimating and evaluating the uncertainty of rating predictions and top-n recommendations in recommender systems. ACM Transactions on Recommender Systems (TORS), 1(2):7:1–7:??, June 2023. ISSN

2770-6699. URL https://dl.acm.org/
doi/10.1145/3584021.

Li:2023:EWG

[8] Xueqi Li, Guoqing Xiao, Yuedan Chen, Zhuo Tang, Wenjun Jiang, and Kenli Li. An explicitly weighted GCN aggregator based on temporal and popularity features for recommendation. ACM Transactions on Recommender Systems (TORS), 1(2):8:1–8:??, June 2023. ISSN

2770-6699. URL https://dl.acm.org/
doi/10.1145/3587272.

Zhou:2023:SSF

[9] Xin Zhou, Aixin Sun, Yong Liu, Jie Zhang, and Chunyan Miao. SelfCF: a simple framework for self-supervised collaborative filtering. ACM Transactions on Recommender Systems (TORS), 1(2):9:1–9:??, June 2023. ISSN 2770-

                                        6699. URL https://dl.acm.org/doi/
                                        10.1145/3591469.

Li:2023:WWP

[10] Ming Li, Mozhdeh Ariannezhad, Andrew Yates, and Maarten De Rijke. Who will purchase this item next? Reverse next period recommendation in grocery shopping. ACM Transactions on Recommender Systems (TORS), 1 (2):10:1–10:??, June 2023. ISSN 2770-

                                        6699. URL https://dl.acm.org/doi/
                                        10.1145/3595384.

Cavenaghi:2023:SSR

[11] Emanuele Cavenaghi, Gabriele Sottocornola, Fabio Stella, and Markus Zanker. A systematic study on reproducibility of reinforcement learning in recommendation systems. ACM Transactions on Recommender Systems (TORS), 1(3):11:1– 11:??, September 2023. ISSN 2770-

                                        6699. URL https://dl.acm.org/doi/
                                        10.1145/3596519.

Tomlinson:2023:TTM

[12] Kiran Tomlinson, Mengting Wan, Cao Lu, Brent Hecht, Jaime Teevan, and Longqi Yang. Targeted training for multiorganization recommendation. ACM Transactions on Recommender Systems (TORS), 1(3):12:1–12:??, September 2023. ISSN 2770-6699. URL https://

                                        dl.acm.org/doi/10.1145/3603508.

Benedict:2023:ISM

[13] Gabriel Bénédict, Daan Odijk, and Maarten de Rijke. Intent-satisfaction modeling: From music to video streaming. ACM Transactions on Recommender Systems (TORS), 1(3):13:1– 13:??, September 2023. ISSN 2770-

6699. URL https://dl.acm.org/doi/
10.1145/3606375.

Rendle:2023:RUI

[14] Steffen Rendle and Li Zhang. On reducing user interaction data for personalization. ACM Transactions on Recommender Systems (TORS), 1(3):14:1–14:??, September 2023. ISSN 2770-6699. URL https:

//dl.acm.org/doi/10.1145/3600097.

Nguyen:2023:TCP

[15] Tung Nguyen and Jeffrey Uhlmann. Tensor completion with provable consistency and fairness guarantees for recommender systems. ACM Transactions on Recommender Systems (TORS), 1(3):15:1– 15:??, September 2023. ISSN 2770-

6699. URL https://dl.acm.org/doi/
10.1145/3604649.